### v1

# Phased Instruction Fine-Tuning for Large Language Models

Abstract
Instruction Fine-Tuning, a method enhancing pre-trained language models’ capabilities from mere next-word prediction to complex instruction following, often employs a one-off training approach on diverse instruction dataset. However, this method may not effectively enhance models’ adherence to instructions due to the simultaneous handling of varying instruction complexities. To address this, we propose a novel phased instruction fine-tuning (Phased IFT) method, grounded in the hypothesis of progressive alignment, which posits that the transition of a pre-trained language model from simple next-word prediction to sophisticated instruction following is a gradual learning process. Specifically, we obtain the score of difficulty for each instruction via GPT-4, stratify the instruction data into subsets of increasing difficulty, and sequentially uptrain on these subsets using the standard supervised loss. Through extensive experiments on the pre-trained models Llama-2 7B/13B, and Mistral-7B using the 52K Alpaca instruction data, we demonstrate that Phased IFT significantly surpasses traditional one-off instruction fine-tuning (One-off IFT) method in win rate, empirically validating the progressive alignment hypothesis. Our findings suggest that Phased IFT offers a simple yet effective pathway for elevating the instruction-following capabilities of pre-trained language models. Models and datasets from our experiments are freely available at https://github.com/xubuvd/PhasedSFT.
Phased Instruction Fine-Tuning for Large Language Models
Wei Pang pangweitf@bupt.cn                        Chuan Zhou Peking University
Xiao-Hua Zhou Peking University                        Xiaojie Wang Beijing University of Posts and Telecommunications
1 Introduction
Figure 1: Win rate comparison between: (the red line) Phased IFT on difficulty-stratified instruction data and (the yellow line) Phased IFT on randomly stratified instruction data, against (the gray dotted line) a baseline of One-off IFT on the whole instruction dataset.
Instruction fine-tuning (IFT) Ouyang et al. (2022); Longpre et al. (2023), involving training on instruction dataset using standard supervised fine-tuning method, aligns pre-trained language models to users’s intent and has been proven as an effective alignment method to enhance their ability to follow instructions. Large language models (LLMs) are pre-trained on raw text data using maximum likelihood estimation, equipping them with the basic ability to predict the next word Zhou et al. (2023); Zhao et al. (2023). However, a gap exists between this ability and following user intentions Thoppilan et al. (2022); Ouyang et al. (2022). To bridge this gap and enable models to complete human end tasks, various instruction fine-tuning strategies have been proposed, including SFT Ouyang et al. (2022), LIMA Zhou et al. (2023), Alpaca Taori et al. (2023); Dubois et al. (2023), Alpagasus Chen et al. (2024), CoT Wei et al. (2022), Superfiltering Li et al. (2024) and Self-instruct Wang et al. (2023b). Among these, SFT and LIMA employ human-written instruction data for fine-tuning, while strategies like Self-instruct, Alpaca, Alpagasus utilize ChatGPT as a teacher to automatically generate extensive instruction datasets. Their training approach, a one-off IFT on the whole instruction data without differentiating the difficulty levels of instructions, lacks efficiency in enhancing the instruction-following capability of pre-trained models.
An instruction sample comprises a triplet of an instruction, an optional input, and an output Wang et al. (2023b). The instruction describes the task, the input serves as an additional context to the instruction, and the output is the answer following the instruction. Diversity Dubois et al. (2023); Wang et al. (2023b); Zhou et al. (2023) is a crucial aspect of a high-quality instruction dataset, indicating that such datasets are typically extensive, encompassing a wide range of tasks and examples representing different levels of instruction-following difficulty. For instance, tasks like mathematical problem-solving, code writing, entity extraction, and copy generation each present varying levels of difficulty. Even within the same category of instructions, the difficulty can vary with the length of the input (refer to Table 1 for details). In the realm of entity extraction, the difficulty differs between extracting place names and identifying product name fragments in e-commerce Li et al. (2023b).
However, previous instruction fine-tuning methods have treated large-scale instruction dataset uniformly, feeding them to the pre-trained model for one-time alignment training without differentiating the role of instructions of varying difficulties in the fine-tuning process. This approach overlooks the nuanced differences in instruction complexities and potentially limits the efficiency of model training.
To address the aforementioned issue, this paper introduces an instruction difficulty scoring mechanism. Specifically, the difficulty of an instruction is defined by the complexity of the instruction itself, the input, and the challenges involved in generating the output. We employ the strongest available models (such as GPT-4 OpenAI (2023)) as a teacher to score the difficulty of each instruction and input, as well as the challenge in generating the output, on a scale from 1 to 5, where higher scores indicate greater difficulty. Further, we plot probability density curve of the instruction difficulty as a heuristic guide Campello et al. (2020), coupled with expert judgment to select thresholds for difficulty scores. This approach enabled us to segment the instruction dataset into multi-stages sub-datasets, forming a sequence with increasing difficulty levels. Building upon this, we further propose the progressive alignment hypothesis: Aligning a pretrained model’s existing ability to predict the next word with the capability to generate content following human intent is a gradual learning process, progressively attaining alignment with human intent.
Based on the hypothesis, we develop a Phased Instruction Fine-Tuning (Phased IFT) method, diverging from the traditional One-off IFT approach. Phased IFT represents an effective sequential uptraining framework, which entails progressively training on sub-instruction datasets of incremental difficulty. Initially, training commences on an easy sub-dataset with the standard supervised loss. Following training completion, the model checkpoint is saved and subsequently utilized to extend supervised training to a marginally more complex sub-dataset. This iterative process persists until the most challenging sub-dataset is addressed.
To assess this method, we engage in extensive experiments using two widely used instruction datasets: Alpaca 52K and its refined variant, Alpaca-clean, in conjunction with three leading open-source pre-trained models: Llama-2 7/13B Touvron et al. (2023) and Mistral-7B Jiang et al. (2023). GPT-4 served as judge to assess the gains of Phased IFT against One-off IFT in win rate. In evaluations across 6 benchmark datasets, substantial enhancements in win rate are recorded for the Phased IFT method. Figure 1 illustrates the win rate as a function of the uptraining stages for phased IFT versus the traditional one-off IFT method. The red line indicates phased IFT on difficulty-staged data, after uptraining in the stage 2, the win rate already surpasses that of the one-off IFT method, and after the stage 3 of uptraining, the win rate significantly increases. The yellow line shows phased IFT on randomly staged data, the win rate after fine-tuning in any stage is lower than that of the traditional one-off IFT method, highlighting the substantial advantage of our proposed phased instruction fine-tuning method. Furthermore, extensive ablation studies validate the hypothesis and the efficacy of the Phased IFT, providing robust empirical support. The main contributions of this paper include:
  • We propose a phased instruction fine-tuning (Phased IFT) method, utilizing GPT-4 for scoring the difficulty of instructions and subsequently dividing the instruction dataset into a sequence of multi-stages sub-datasets with increasing difficulty levels. This approach employs uptraining on the sequence of multi-stages sub-datasets.
  • We introduce the progressive alignment hypothesis, suggesting that aligning the generative abilities of pre-trained models with human instructions is a process of incremental training, rather than achievable through one-off fine-tuning. This hypothesis is substantiated through extensive experimental validation.
  • A series of experiments demonstrate that our Phased IFT method is more effective than One-off IFT approach. In terms of win rate across six benchmark, the Phased IFT shows significant improvements.
2 Related Work
Open-source Instruction-tuned Models Recent studies have shown that fine-tuning Large Language Models (LLMs) with open-source instruction datasets significantly boosts their adherence to human instructions. The 65B parameter LIMA Zhou et al. (2023) model, after fine-tuning with 1,000 human-generated instructions, outperformed GPT-4 in 43% of cases, Bard in 58%, and DaVinci003 in 65%, showcasing its superior instruction-following capability. Similarly, the Alpaca 7B model Taori et al. (2023), fine-tuned on 52K instructions from text-davinci-003, matched its performance, demonstrating the effectiveness of large-scale, diverse dataset fine-tuning. Advancing further, Alpagasus Chen et al. (2024) selects a subset of 9k high-quality instructions from the Alpaca 52K and is fine-tuned on the subset, significantly surpassing the original Alpaca’s Taori et al. (2023) performance across four human-instruction test sets, e.g., Self-Instruct Wang et al. (2023b), WizardLM Xu et al. (2024), Koala Geng et al. (2023) and Vicuna Chiang et al. (2023), highlighting the benefits of precise instruction selection and targeted fine-tuning for enhancing LLM capabilities. Specifically, Alpagasus prompts ChatGPT to score each instruction on a scale of 0 to 5, resulting in a predominant score range of 3 to 4, comprising roughly 78.9% of its total. Then Alpagasus selectively utilizes 9,229 instructions rated above 4.5, establishing a high-quality dataset for fine-tuning.
Open-source Instruction Dataset The Alpaca dataset Taori et al. (2023), designed to improve the instruction-following capabilities of language models and comprising 52K instructions, was generated using OpenAI’s text-davinci-003. This refined the Self-Instruct Wang et al. (2023b) framework, resulting in a dataset with significantly greater diversity than its Self-Instruct counterpart. Alpaca-cleaned yahma (2023) is a cleaned version of the original Alpaca dataset, identified several issues, such as hallucinations, empty outputs and wrong answers, thereby elevating the overall quality of the data. Additionally, LIMA Zhou et al. (2023) released a meticulously curated dataset comprising 1000 instructions, incorporating 750 popular queries and responses from community forums like Stack Exchange and wikiHow, alongside 250 expertly crafted entries. This compilation emphasizes the quality and diversity of instructions.
Instruction Fine-tuning Strategy The prevalent method of instruction fine-tuning, employed in models such as LIMA, Alpaca, Alpagasus and Superfiltering Li et al. (2023a, 2024), involves a one-off instruction fine-tuning (One-off IFT) on the whole instruction dataset characterized by their diversity and high quality. This approach, while emphasizing dataset quality and diversity Wang et al. (2023b); Li et al. (2023a), neglects the inherent complexity of the instruction sets. Consequently, this one-off IFT fails to adequately equip models with the nuanced capability to comprehend and execute a wide array of instructions. Differs from the one-off IFT approach, our proposed phased IFT method initially segments the instruction dataset into multiple sub-datasets, arranged in a sequence from simple to complex in terms of instruction difficulty. Continuous uptraining is then executed across this sequence.
3 Method
Figure 2: Overview of the proposed Phased Instruction Fine-Tuning (Phased IFT).
Figure 2 presents the overall pipeline of our proposed Phased Instruction Fine-Tuning (Phased IFT), encompassing three key components: Instruction Difficulty Calculating (IDC), Instruction Dataset Segmenting (IDS), and Uptraining. Where, IDC prompts GPT-4 to allocate a difficulty rating to each instruction. Following this, IDS converts the instructions’s score into a cumulative probability distribution, which facilitates the division of the dataset into subsets categorized by difficulty level, and segments the whole instruction data into several stages. Finally, across this stages, Uptraining is performed, beginning with those classified as less challenging. Detailed elaborations on the specifics of each component are provided in the subsequent sections of this chapter.
Figure 3: A prompt to ChatGPT-4 for scoring instruction difficulty.
Instruction Difficulty Calculating (IDC) Following Alpagasus’s prompt Chen et al. (2024), we also design a GPT-4-based scoring prompt to assess the difficulty of each triplet of (instruction, input, output) within the 52K Alpaca dataset, using a scale of 1 to 5, with higher scores indicating greater difficulty. This prompt is detailed in Figure 3, where each triplet is evaluated in two dimensions. Firstly, we consider the intrinsic difficulty of the instruction itself, reflecting the complexity inherent to the task described. Secondly, we evaluate the challenge involved in transforming the given instruction to the expected output, thereby estimating the difficulty from instruction to output. This dual-dimensional scoring approach allows for a nuanced understanding of task complexity within the dataset.
The difficulty score histogram of the 52K Alpaca dataset is illustrated as shown in the left of Figure 2. Another interesting finding is that a majority of instructions in the 52K Alpaca are categorized as low difficulty, with approximately 63.5% scoring below 1.5. In contrast, within the Alpaca-clean dataset, instructions with a difficulty score lower than 1.5 account for 52.3%; in the LIMA dataset, the proportion of instructions with a difficulty score below 1.5 is merely 9.9%, while those scoring above 4.5 account for a significant 58.3%. Details of their comparison can be found in Appendix A.1. This highlights the feasibility of employing GPT-4 to assess the difficulty level of instructions.
Instruction Dataset Segmenting (IDS)
Our goal is to divide the instruction dataset into multi-stages sub-datasets based on difficulty scores, while ensuring that the number of samples in each sub-dataset remains relatively balanced. To obtain heuristic information from the scored, discrete dataset, we plot Gaussian kernel density curve on the scored dataset. This curve represents the cumulative probability density as the difficulty score increases from the minimum to the current score, as depicted in the middle part of Figure 2. Utilizing human experience, we select threshold scores, such as the 1.5 and 3.5 marked by blue vertical lines. These thresholds segment the whole 52K Alpaca instruction dataset into three stages sub-datasets.
The first stage (stage 1), covering difficulty scores from [1, 1.5), encompasses 30.3k samples; the second stage (stage 2), spanning scores from [1.5, 3.5), accounts for 14.7k samples; and the third stage (stage 3), with scores within [3.5, 5], comprises 6.8k samples. These three stages sub-datasets form a sequence of sub-datasets with increasing difficulty scores, which are used for uptraining in the next phase.
Uptraining In the post-IDS phase, uptraining is performed across three sequential stages, initiating with fine-tuning at the first stage and progressively advancing through the second and third stages. Throughout this process, hyperparameters such as learning rate, batch size, epoch count, and warmup ratio are consistently applied. A standard supervised loss function is used in the uptraining. Specifically, the triplet of (instruction, input, output) is merged into a text string for training, wherein the model predicts each token. Nonetheless, loss calculation is confined to the output part, with loss from the instruction and input parts masked.
4 Experiments
This paper seeks to validate the progressive alignment hypothesis, examining the relative efficacy of two instruction alignment training methodologies: One-off Instruction Fine-Tuning (One-off IFT) versus our proposed Phased Instruction Fine-Tuning (Phased IFT), through three distinct experiments. Firstly, Experiment 1 evidences Phased IFT’s superior performance over One-off IFT across two instruction data, establishing Phased IFT’s effectiveness. Secondly, Experiment 2 assesses the merit of stratifying instruction datasets by difficulty versus arbitrary segmentation, aiming to attribute Phased IFT’s success to difficulty-stratified instruction data categorization rather than to the phased training approach per se. Finally, Experiment 3, conducting a full permutation analysis of Phased IFT, seeks to demonstrate that the method’s success is specifically due to uptraining in a sequence of increasing difficulty, with alternative sequencing failing to produce comparable outcomes.
Datasets and Evaluation We conducted extensive experiments on the standard Alpaca 52K instruction data and its refined version Alpaca-cleaned, with the following three recent LLMs: Llama-2 7/13B and Mistral-7B respectively, evaluating them across 6 benchmark datasets, including Self-Instruct of size 252, WizardLM of size 218, Koala of size 180, Vicuna of size 80, OASST of size 188 Köpf et al. (2023), and Anthropic Bai et al. (2022) of size 129.
Following evaluation from Alpagasus Chen et al. (2024), we prompt GPT-4 to judge Win-Tie-Lose and compute a win rate. Under identical instruction and input, GPT-4 as an auto-grader Zheng et al. (2023); Chia et al. (2023) to score the quality of output from both One-off IFT and Phased IFT methods, on a scale of 1 to 10 (See prompt in Appendix A.2). For fairness, we score each instruction twice, i.e., placing Phased IFT’s output before/after the One-off IFT’s output. An instruction’s overall outcome is defined by:
  • Win: win twice or win once and tie once;
  • Tie: tie twice or win once and loses once;
  • Lose: loses twice or loses once and tie once.
The win rate calculation formula is as follows:
$\mathrm{WinRate}=\frac{\#win+\#tie/2}{\#win+\#tie+\#lose}-0.5,$
where # means the corresponding number of samples.
Table 1: Detailed statistics in each stage of Alpaca and Alpaca-cleaned instruction data, including the number of samples (K), average difficulty per sample, and average token length per sample.
Data Number (K) Avg. Difficulty Avg. Token Length
stage1 stage2 stage3 stage1 stage2 stage3 stage1 stage2 stage3
Alpaca 30.3 14.7 6.8 1.00 2.21 4.25 63.16 106.83 134.40
Alpaca-cleaned 23.9 16.4 11.4 1.00 2.22 4.25 113.42 200.99 301.60
Stratifying Alpaca by Difficulty Score
Figure 4: Difficulty score histogram of Alpaca-cleaned data and its cumulative probability density.
The instruction difficulty within the Alpaca and Alpaca-cleaned data are quantitatively assessed by GPT-4, assigning scores from 1 to 5, with higher score denoting increased complexity. Based on the density curve (e.g., Figure 2 for Alpaca and Figure 4 for Alpaca-cleaned), we adopt two threshold values, namely 1.5 and 3.5 for both Alpaca and Alpaca-cleaned, and segment them into three stages sub-datasets, respectively. We refer to as the Alpaca-3-stages and Alpaca-cleaned-3-stages.
Additionally, to conduct ablation experiments, we also construct a randomly sampled 3-stages data, with each stage containing the same number of instructions as the difficulty-segmented subset. This is referred to as the Alpaca-rand-3-stages and Alpaca-cleaned-rand-3-stages data. Detailed statistics, as delineated in Table 1.
Implementations We utilized Huggingface’s Trainer with DeepSpeed’s Zero3 for data-parallel, full-parameter fine-tuning in our training regimen. Consistently across experiments, a learning rate of 5e-6 and a duration of two training epochs were applied. We set a per-device batch size of 16 and a gradient accumulation steps of 1, a weight decay of 0.1, a warmup ratio of 0.1, and a cosine annealing learning rate scheduler.
4.1 Experiment 1: Comparison win rate of Phased IFT and One-off IFT
Table 2: Comparisons of win rates between Phased IFT and One-off IFT. Phased IFT involves up-training on Alpaca-3-stages and Alpaca-cleaned-3-stages, while One-off IFT is fine-tuned solely on the original Alpaca and Alpaca-cleaned data. The changes in win rates are represented by percentage differences, where a "+" indicates an increase and "-" denotes a decrease in percentage.
Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Data
Llama-2 7B +5.75 +10.32 +4.17 +11.25 +8.78 +12.40 Alpaca
Mistral-7B -0.20 -2.52 +6.67 +5.00 +3.19 +3.10
Llama 2-13B +4.37 +6.88 +10.56 +11.25 +10.11 +13.18
Llama 2-7B +3.77 +4.13 +4.17 +3.12 - - Alpaca-cleaned
Llama 2-13B +2.78 +3.21 +1.39 +6.88 +7.45 +9.30
Table 2 demonstrates that the Phased IFT method yields superior performance across a wide range of evaluation metrics compared to the One-off IFT method, with the sole exceptions being the Mistral-7B model’s lower win rates on the Self-Instruct and WizardLM benchmarks. Specifically, the Llama-2 7B model, when fine-tuned with the Phased IFT method on the Alpaca-3-stages instruction set, experienced a win rate increase ranging from a minimum of +4.17 absolute points (on Koala) to a maximum of +12.40 absolute points (on Anthropic). Similarly, the Llama-2 13B model’s win rate improved by at least +4.37 points (on Self-Instruct) to as much as +13.18 absolute points (on Anthropic). Notably, the Mistral-7B model saw a win rate increase of +6.67 absolute points on Koala and +3.10 points on Anthropic, but it exhibited a decrease in win rates of -0.20 and -2.52 absolute points on Self-Instruct and WizardLM, respectively.
On the Alpaca-cleaned instruction set, the win rate of the Llama 2-7B model fine-tuned via Phased IFT improved by at least +3.12 absolute points (on Vicuna) and up to +4.17 points (on Koala), while the Llama-2 13B model saw increases ranging from +1.39 absolute points (on Koala) to +9.30 points on Anthropic.
These results from Experiment 1 affirm the efficacy of the Phased IFT method for instruction-based fine-tuning of pre-trained LLMs, leading to notable performance enhancements on Llama-2 7B/13B and Mistral-7B models.
4.2 Experiment 2: Comparison win rate of difficulty-stratified 3-stages and randomly sampled 3-stages
Table 3: Comparison of win rates between Phased IFT and One-off IFT methods is outlined in three rows. The first row (Row #1) assesses Phased IFT on Alpaca-rand-3-stages against One-off IFT using the original Alpaca 52K data. In the second row (Row #2), the comparison is made between Phased IFT on Alpaca-cleaned-rand-3-stages and One-off IFT on the original Alpaca-cleaned data. The third row (Row #3) evaluates the win rates of Phased IFT on Alpaca-3-stages in contrast to Phased IFT on Alpaca-rand-3-stages. Each entry corresponds to Table 2.
Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Row
Llama-2 7B -1.19 -2.75 +0.56 +0.62 -0.66 +0.75 #1
Llama-2 13B -3.97 -3.67 -3.89 -3.75 -2.93 -3.49
Llama-2 7B -2.78 -2.52 -5.00 -1.88 - - #2
Llama-2 7B +9.13 +12.16 +13.61 +10.62 +5.85 +6.59 #3
Llama-2 13B +4.76 +7.57 +1.94 +10.00 +10.37 +10.85
Table 3 provides insight into the effectiveness of our Phased IFT method for fine-tuning pre-trained models Llama-2 7B and 13B on randomly sampled 3-stages instruction data. Notably, the Phased IFT method, in comparison to the One-off IFT approach, exhibited a decline in win rates across several test sets. Specifically, as listed in the Row #1 fine-tuned with Phased IFT on Alpaca-rand-3-stages, the win rate of the Llama-2 13B model decreased significantly compared to One-off IFT on the original Alpaca 52K data, with reductions ranging from -2.93 absolute points on Anthropic to -3.97 points on Self-Instruct.
The Row #2 of Table 3 which focuses on the application of Phased IFT for fine-tuning Llama-2 7B on Alpaca-cleaned-rand-3-stages, shows a pronounced decrease in win rate compared to the One-off IFT on the original Alpaca-cleaned data, with declines spanning from -1.88 absolute points on Vicuna to -5.00 absolute points on Koala.
Contrastingly, Row #3 of Table 3 demonstrates that applying Phased IFT for uptraining the Llama-2 7B and 13B models on Alpaca-3-stages, as opposed to on Alpaca-rand-3-stages, yielded significant improvements in win rate. Specifically, the Llama-2 7B model saw increases from +5.85 absolute points on OASST to +13.61 on Koala, while the Llama-2 13B model’s performance improved by +1.94 points on Koala and up to +10.85 points on Anthropic.
These results indicate that uptraining with randomly selected 3-stages instruction data does not inherently enhance performance and can lead to reductions across most test sets. This contrasts with experiment 1 where Phased IFT applied to difficulty-stratified 3-stages instruction sets, matched for instruction count and compared against the same baseline, yielded substantial performance gains. This highlights two key insights: firstly, the improvements observed are not solely attributable to the 3-stages uptraining process; secondly, categorizing instruction data by difficulty level provides a structured fine-tuning approach, enabling pre-trained models to progressively adapt to more complex tasks.
Table 4: Win rate of all permutations of uptraining of Llama-2 7B/13B on Alpaca-3-stages instruction data, compared to One-off IFT on the original Alpaca data. Each entry is the same as in Table 2.
Permutations Self-Instruct WizardLM Koala Vicuna OASST Anthropic Base Models
1-2-3 +5.75 +10.32 +4.17 +11.25 +8.78 +12.40 7B
2-1-3 +0.40 +2.98 +0.56 +11.25 +6.65 +8.14
3-1-2 +3.57 +1.15 -3.33 +12.50 +1.06 +5.43
1-3-2 -0.79 -2.52 -0.83 +2.50 -2.39 +5.43
2-3-1 -5.16 -9.40 -13.61 -8.12 -5.32 -5.04
3-2-1 -4.76 -5.96 -9.17 -6.25 -9.04 -1.16
1-2-3 +4.37 +6.88 +10.56 +11.25 +10.11 +13.18 13B
2-1-3 +6.15 +7.11 +10.28 +5.62 +5.85 +8.53
3-1-2 +0.79 +1.38 +4.44 -6.88 +2.39 +2.33
1-3-2 +1.39 +3.90 +5.00 +1.25 +1.06 +6.98
2-3-1 -3.37 -6.65 -1.11 -7.50 -11.70 -10.08
3-2-1 -5.16 -5.28 -5.00 -11.88 -12.77 -7.36
4.3 Experiment 3: Comparison win rate of uptraining on Alpaca-3-stages in all permutations
To investigate the impact of varying the sequence of Alpaca-3-stages on the performance of the Llama-2 7B/13B models with clarity and precision, the following structure is proposed. This structure presents the uptraining order across the three stages, denoted as Stage 1 (1), Stage 2 (2), and Stage 3 (3), to illustrate the sequence in which the model was uptrained on different segments of the instruction data:
  • 1-2-3: This order indicates that the model was first uptrained on data from Stage 1, followed by Stage 2, and finally Stage 3.
  • 2-1-3: The model was initially uptrained on Stage 2 data, then on Stage 1, and finally with Stage 3.
  • 3-1-2: The uptraining commenced with Stage 3 data, proceeded to Stage 1, and ended with Stage 2.
  • 1-3-2: The sequence started with Stage 1, moved to Stage 3, and finished with Stage 2.
  • 2-3-1: Uptraining began with Stage 2 data, followed by Stage 3, and then Stage 1.
  • 3-2-1: This order shows that the model was first uptrained on Stage 3 data, then Stage 2, and finally Stage 1.
Table 4 demonstrates the performance enhancements achieved by fine-tuning the pretrained Llama-2 7B/13B models through the Phased IFT method across 6 permutations of these stages, in contrast to the One-off IFT method applied directly on the original Alpaca data. The results distinctly reveal that the Llama-2 7B/13B models achieve maximal win-rate improvements when sequentially fine-tune in the 1-2-3 order using Phased IFT, substantially outperforming other sequential arrangements.
The Llama-2 7B model, after undergoing Phased IFT in a 1-2-3 sequence, demonstrates significant win-rate enhancements of +5.75, +10.32, +4.17, +11.25, +8.78, and +12.40 absolute points across Self-Instruct, WizardLM, Koala, Vicuna, OASST, and Anthropic benchmarks, respectively. This outperformance is substantial when compared to the model fine-tuned via the One-off IFT method on the 52K Alpaca data, as well as models fine-tuned in alternative sequences. Notably, the 1-2-3 and 2-1-3 training sequences yield markedly superior win rate than the 3-1-2 and 1-3-2 sequences, which in turn outperform the 2-3-1 and 3-2-1 sequences. The latter sequences exhibit win rate significantly lower than those achieved through One-off IFT.
Similarly, the Llama-2 13B model, fine-tuned in the 1-2-3 sequence using Phased IFT, show win-rate improvements of +4.37, +6.88, +10.56, +11.25, +10.11, and +13.18 absolute points on the same benchmarks, respectively, again surpassing the performance of models fine-tuned through other sequences or the One-off IFT method. The training sequences of 1-2-3 and 2-1-3 consistently resulted in higher win rate compared to 3-1-2 and 1-3-2, which outperform 2-3-1 and 3-2-1 sequences. The win rate for the 2-3-1 and 3-2-1 sequences are notably inferior to those achieved with One-off IFT.
These findings compellingly support the hypothesis that uptraining in order of increasing difficulty presents significant positive gains, whereas reverse order training yields negligible benefits. This underpins our progressive alignment hypothesis, positing that the process of aligning pretrained models to human intent through supervised instruction-based fine-tuning is inherently gradual. Beginning with basic and straightforward instructions before progressively tackling more challenging ones ensures a systematic alignment with end tasks. This approach not only validates the utility of constructing diverse and difficulty-tiered instruction data but also highlights the strategic importance of adopting a progressively challenging fine-tuning regimen to enhance model performance and alignment with human intent in task completion.
To summarize, Experiment 1 validates the superiority of phased instruction fine-tuning (Phased IFT) over one-off IFT, evidenced by a notable enhancement in win rates. Experiment 2 underscores the significance of stratifying instruction dataset by difficulty score, yielding a pronounced improvement in win rate against random sampling for uptraining fine-tuning, as depicted in Row #3 of Table 3, with the former method achieving a significant uplift in win rate. Experiment 3 demonstrates that sequential uptraining fine-tuning by ascending difficulty represents the optimal phased training approach, concurrently corroborating the progressive alignment hypothesis. Collectively, these experiments show the efficacy of the progressive alignment hypothesis and Phased IFT.
5 Conclusion
The paper introduces the progressive alignment hypothesis, suggesting that the transition of a pretrained model’s next-word prediction capability into an ability to follow human intentions is a sequential learning process. This process entails progressively aligning with human intent, advancing from simpler to more complex tasks. In light of this hypothesis, we propose a phased instruction fine-tuning (Phased IFT) approach. Utilizing GPT-4 to score the difficulty of instructions, we categorize the instruction dataset into different stages according to the difficulty score density distribution. Diverging from prior methods that apply a one-off instruction fine-tuning (One-off IFT) on the whole instruction data, our approach entails staged, supervised uptraining, addressing each stage in ascending order of difficulty. Through extensive experiments, we have validated the progressive alignment hypothesis and the efficacy of Phased IFT, demonstrating its utility in developing diverse instruction dataset and enhancing instruction fine-tuning.
6 Broader Impact
This paper discusses a fundamental issue in Large Language Models (LLMs): how to align LLMs more efficiently. The work introduces a simple and effective method for alignment, which involves analyzing the difficulty of instruction dataset and structuring the dataset into several stages, followed by supervised fine-tuning through uptraining. We conduct three experiments, with an additional experiment detailed in Appendix A.3, to actively validate the existence of the progressive alignment hypothesis. If the progressive alignment hypothesis is a pervasive phenomenon, this work is poised to have a widespread impact on the field of LLMs.
The datasets employed in this work, namely Alpaca 52K and Alpaca-cleaned, are publicly available, and our training framework leverages open-source tools, utilizing Hugging Face’s Trainer and DeepSpeed. Furthermore, we have purchased access to OpenAI’s API to employ GPT-4 for assessing the difficulty of instruction datasets and for evaluating the quality of generated responses. We will make all our code and data openly available, including training code, GPT-4 difficulty scoring code, win-rate calculation code, and the difficulty-stratified Alpaca and Alpaca-cleaned instruction datasets.
7 Limitations
Here, we discuss some limitations of our study to inspire future research in this direction.
Automatic instruction dataset segmentation This work employs heuristic information from the cumulative density curve of difficulty scores, supplemented by human judgment for threshold selection. A significant challenge lies in developing algorithms capable of autonomously identifying threshold points from this curve, eliminating the reliance on human intuition and ensuring balanced data size across multi-stages sub-datasets.
Validation on larger-scale LLMs Our empirical analysis is confined to pre-trained models with capacities of 7B and 13B, lacking empirical validation on larger-scale models, such as 65B or 70B. This limitation suggests a necessity for extending validation efforts to encompass a wider range of model sizes to enhance the generalizability of our findings.
Acknowledgements
We thank the reviewers for their comments and suggestions. This paper was partially supported by the National Natural Science Foundation of China (NSFC 62076032), Huawei Noah’s Ark Lab, MoECMCC “Artificial Intelligence” Project (No. MCM20190701), Beijing Natural Science Foundation (Grant No. 4204100), and BUPT Excellent Ph.D. Students Foundation (No. CX2020309).
Appendix A Appendix
A.1 Histogram and density curve for Alpaca, Alpaca-cleaned and LIMA
Figure 5: Histogram and cumulative probability density of difficulty scores for Alpaca 52K dataset.
Figure 6: Histogram and cumulative probability density of difficulty scores for Alpaca-cleaned 52K dataset.
Figure 7: Histogram and cumulative probability density of difficulty scores for LIMA 1000 dataset.
We conducted a comparative analysis on the instruction difficulty of three datasets—Alpaca, Alpaca-cleaned, and LIMA—employing GPT-4 for scoring. Figures 5, 6 to 7 illustrate their respective score distributions and probability density functions. Alpaca, generated by the Self-Instruct framework, consists of 52K triplets (instruction, input, output). Alpaca-cleaned represents an enhanced version of Alpaca, refined to minimize errors such as incorrect responses and hallucinations. LIMA, by contrast, is composed of 1,000 manually curated instructions. Analysis of score histograms across these figures indicates a predominance of higher scores in LIMA, comparatively fewer in Alpaca, with Alpaca-cleaned exhibiting overall improved scoring relative to Alpaca. These findings corroborate our qualitative assessment of the datasets and affirm the utility of GPT-4 in evaluating instruction difficulty.
The following list includes seven examples of instructions scored by GPT-4, with respective scores of 1, 1.5, 2.6, 3.25, 4.5, 4.75, and 5. The scoring prompt is provided in Figure 3.
  • Score: 1.0, Instruction: Given the following definition, build a sentence using the word. Input: To send away abruptly or angrily. Output: He angrily dismissed me without even giving me an explanation.
  • Score: 1.5, Instruction: Generate two points that could be used to extend the following statement. Input: Internet has become an important way to access knowledge. Output: - It has also become a key tool for collaboration and communication. - It has provided a platform for the sharing of ideas and greater access to global markets.
  • Score: 2.6, Instruction: Propose two appropriate and different methods to solve the given problem. Input: Predicting whether an email is a spam or not. Output: Two possible methods for predicting whether an email is a spam or not are supervised machine learning models such as artificial neural networks and Naive Bayes classifiers, or an unsupervised machine learning method such as clustering and anomaly detection.
  • Score: 3.25, Instruction: Edit this passage to make it more engaging to readers. Input: We went for a walk in the park and saw a lot of trees. Output: As we strolled through the crisp park, we were treated to sights of lush green trees rustling in the wind, stretching for miles with their towering dignity.
  • Score: 4.5, Instruction: What is a 9-digit number that is divisible by 7? Input:. Output: 105,714,285 is a 9-digit number that is divisible by 7.
  • Score: 4.75, Instruction: Generate a sentence that has an implied metaphor. Input:. Output: My love for her was a warm summer breeze, picking me up and carrying me away.
  • Score: 5, Instruction: Design a logo for a restaurant called ’Taste of Home’.Input:. Output: A font-based logo in a handwritten script style, with the words ’Taste of Home’ in an orange color for a warm and inviting feeling. The logo should also contain an image of a fork and spoon in the same color.
A.2 ChatGPT-4 for evaluating win rate
Figure 8: A prompt to ChatGPT-4 for rating model’s responses with a score between 1 and 10.
Figure 8 details the prompt used for scoring and comparing the quality of output generated by two models using GPT-4. To mitigate any potential positional bias Wang et al. (2023a); Chen et al. (2024) from GPT-4, we conduct two scoring sessions. Specifically, we provide GPT-4 with inputs in two different orders: (instruction, input, output-1 of model 1, output-2 of model 2) and (instruction, input, output-2 of model 2, output-1 of model 1) for separate evaluations.
A.3 Experiment 4: Comparison win rate of IFT only on the Alpaca stage3 data vs. IFT on all the 52k Alpaca
Table 5: Win rate comparison of One-off IFT only on the third stage selected from Alpaca-3-stages against One-off IFT on the original Alpaca 52K data, using Llama-2 7B model. Each entry is the same as in Table 2.
Self-Instruct WizardLM Koala Vicuna OASST Anthropic
+0.79 +6.65 +2.50 +5.00 +3.81 +4.75
We compared the win rates of IFT conducted solely on the third stage, which includes 6.8K samples, against IFT performed on the entire 52K Alpaca data, as shown in Table 5. It was observed that the win rate for IFT on just stage 3 exceeds that of the full 52K dataset, aligning with conclusions drawn in Alpagasus Chen et al. (2024) that the quality of the instruction dataset is crucial. However, a comparison with Experiment 3 in Table 4 reveals that the win rate for IFT solely on stage 3 is significantly lower than when Phased IFT is conducted across all three stages, roughly only half of the latter’s. This underscores the existence of the progressive alignment hypothesis and validates the effectiveness of the Phased IFT approach.

### v2

@@
 Abstract
-Instruction Fine-Tuning, a method enhancing pre-trained language models’ capabilities from mere next-word prediction to complex instruction following, often employs a one-off training approach on diverse instruction dataset.
-However, this method may not effectively enhance models’ adherence to instructions due to the simultaneous handling of varying instruction complexities.
-To address this, we propose a novel phased instruction fine-tuning (Phased IFT) method, grounded in the hypothesis of progressive alignment, which posits that the transition of a pre-trained language model from simple next-word prediction to sophisticated instruction following is a gradual learning process.
-Specifically, we obtain the score of difficulty for each instruction via GPT-4, stratify the instruction data into subsets of increasing difficulty, and sequentially uptrain on these subsets using the standard supervised loss.
-Through extensive experiments on the pre-trained models Llama-2 7B/13B, and Mistral-7B using the 52K Alpaca instruction data, we demonstrate that Phased IFT significantly surpasses traditional one-off instruction fine-tuning (One-off IFT) method in win rate, empirically validating the progressive alignment hypothesis.
-Our findings suggest that Phased IFT offers a simple yet effective pathway for elevating the instruction-following capabilities of pre-trained language models.
-Models and datasets from our experiments are freely available at https://github.com/xubuvd/PhasedSFT.
+Instruction Fine-Tuning enhances pre-trained language models from basic next-word prediction to complex instruction-following.
+However, existing One-off Instruction Fine-Tuning (One-off IFT) method, applied on a diverse instruction, may not effectively boost models’ adherence to instructions due to the simultaneous handling of varying instruction complexities.
+To improve this, Phased Instruction Fine-Tuning (Phased IFT) is proposed, based on the idea that learning to follow instructions is a gradual process.
+It assesses instruction difficulty using GPT-4, divides the instruction data into subsets of increasing difficulty, and uptrains the model sequentially on these subsets.
+Experiments with Llama-2 7B/13B/70B, Llama3 8/70B and Mistral-7B models using Alpaca data show that Phased IFT significantly outperforms One-off IFT, supporting the progressive alignment hypothesis and providing a simple and efficient way to enhance large language models.
+Codes and datasets from our experiments are freely available at https://github.com/xubuvd/PhasedSFT.
 Phased Instruction Fine-Tuning for Large Language Models
@@
 Xiao-Hua Zhou Peking University                        Xiaojie Wang Beijing University of Posts and Telecommunications
+[…]
 1 Introduction
-Figure 1: Win rate comparison between: (the red line) Phased IFT on difficulty-stratified instruction data and (the yellow line) Phased IFT on randomly stratified instruction data, against (the gray dotted line) a baseline of One-off IFT on the whole instruction dataset.
+Figure 1: In the context of increasing difficulty multi-stage sub-datasets, the trend of win rate growth for uptraining (Phased IFT) compared to One-off on the original dataset is observed.
+The gray horizontal line represents the performance baseline of One-off.
 Instruction fine-tuning (IFT) Ouyang et al.
@@
 This approach enabled us to segment the instruction dataset into multi-stages sub-datasets, forming a sequence with increasing difficulty levels.
-Building upon this, we further propose the progressive alignment hypothesis: Aligning a pretrained model’s existing ability to predict the next word with the capability to generate content following human intent is a gradual learning process, progressively attaining alignment with human intent.
+Building upon this, we further propose the Progressive Alignment Hypothesis: Aligning a pre-trained model’s existing ability to predict the next word with the capability to generate content following human intent is a gradual learning process, progressively attaining alignment with human intent.
 Based on the hypothesis, we develop a Phased Instruction Fine-Tuning (Phased IFT) method, diverging from the traditional One-off IFT approach.
@@
 This iterative process persists until the most challenging sub-dataset is addressed.
-To assess this method, we engage in extensive experiments using two widely used instruction datasets: Alpaca 52K and its refined variant, Alpaca-clean, in conjunction with three leading open-source pre-trained models: Llama-2 7/13B Touvron et al.
-(2023) and Mistral-7B Jiang et al.
-(2023).
-GPT-4 served as judge to assess the gains of Phased IFT against One-off IFT in win rate.
-In evaluations across 6 benchmark datasets, substantial enhancements in win rate are recorded for the Phased IFT method.
-Figure 1 illustrates the win rate as a function of the uptraining stages for phased IFT versus the traditional one-off IFT method.
-The red line indicates phased IFT on difficulty-staged data, after uptraining in the stage 2, the win rate already surpasses that of the one-off IFT method, and after the stage 3 of uptraining, the win rate significantly increases.
-The yellow line shows phased IFT on randomly staged data, the win rate after fine-tuning in any stage is lower than that of the traditional one-off IFT method, highlighting the substantial advantage of our proposed phased instruction fine-tuning method.
-Furthermore, extensive ablation studies validate the hypothesis and the efficacy of the Phased IFT, providing robust empirical support.
+We conducted extensive experiments using two leading open-source pre-trained models, Llama2 7/13/70B Touvron et al.
+(2023), Llama3 8/70B, and Mistral 7B Jiang et al.
+(2023), on two widely-used instruction datasets, Alpaca and its refined version AlpacaClean.
+To evaluate the training effectiveness, we utilized 6 benchmarks, with GPT-4 serving as the judge to measure the performance improvements of Phased IFT over One-off IFT in terms of win rate.
+Specially, Figure 1 illustrates the process wherein Phased IFT’s win rate progressively surpasses that of One-off IFT on the three-stage difficulty-increasing sub-datasets.
+It is evident that the three Llama2 models and the two Llama3 models exhibit a consistent trend of increasing win rates with the progression of uptraining on multi-stage sub-datasets, indicating the effectiveness of the progressive alignment hypothesis.
+Furthermore, we conducted three sets of experiments and an ablation study to confirm the effectiveness of Phased IFT.
 The main contributions of this paper include:
@@
 This approach employs uptraining on the sequence of multi-stages sub-datasets.
-  • We introduce the progressive alignment hypothesis, suggesting that aligning the generative abilities of pre-trained models with human instructions is a process of incremental training, rather than achievable through one-off fine-tuning.
-This hypothesis is substantiated through extensive experimental validation.
-  • A series of experiments demonstrate that our Phased IFT method is more effective than One-off IFT approach.
-In terms of win rate across six benchmark, the Phased IFT shows significant improvements.
+  • We introduce the Progressive Alignment Hypothesis, which posits that aligning the generative capabilities of pre-trained models with human intent is a gradual process, rather than being achieved through a one-time fine-tuning on an instruction dataset.
+This hypothesis is supported by extensive experimental validation.
+  • Extensive experiments have demonstrated that our proposed Phased IFT is more effective than One-off IFT.
+It has significantly outperformed One-off IFT across 6 benchmarks.
+Additionally, a series of ablation experiments have validated the effectiveness of the Progressive Alignment Hypothesis.
 2 Related Work
-Open-source Instruction-tuned Models Recent studies have shown that fine-tuning Large Language Models (LLMs) with open-source instruction datasets significantly boosts their adherence to human instructions.
-The 65B parameter LIMA Zhou et al.
-(2023) model, after fine-tuning with 1,000 human-generated instructions, outperformed GPT-4 in 43% of cases, Bard in 58%, and DaVinci003 in 65%, showcasing its superior instruction-following capability.
-Similarly, the Alpaca 7B model Taori et al.
-(2023), fine-tuned on 52K instructions from text-davinci-003, matched its performance, demonstrating the effectiveness of large-scale, diverse dataset fine-tuning.
-Advancing further, Alpagasus Chen et al.
-(2024) selects a subset of 9k high-quality instructions from the Alpaca 52K and is fine-tuned on the subset, significantly surpassing the original Alpaca’s Taori et al.
-(2023) performance across four human-instruction test sets, e.g., Self-Instruct Wang et al.
-(2023b), WizardLM Xu et al.
-(2024), Koala Geng et al.
-(2023) and Vicuna Chiang et al.
-(2023), highlighting the benefits of precise instruction selection and targeted fine-tuning for enhancing LLM capabilities.
-Specifically, Alpagasus prompts ChatGPT to score each instruction on a scale of 0 to 5, resulting in a predominant score range of 3 to 4, comprising roughly 78.9% of its total.
-Then Alpagasus selectively utilizes 9,229 instructions rated above 4.5, establishing a high-quality dataset for fine-tuning.
 Open-source Instruction Dataset The Alpaca dataset Taori et al.
-(2023), designed to improve the instruction-following capabilities of language models and comprising 52K instructions, was generated using OpenAI’s text-davinci-003.
-This refined the Self-Instruct Wang et al.
-(2023b) framework, resulting in a dataset with significantly greater diversity than its Self-Instruct counterpart.
-Alpaca-cleaned yahma (2023) is a cleaned version of the original Alpaca dataset, identified several issues, such as hallucinations, empty outputs and wrong answers, thereby elevating the overall quality of the data.
-Additionally, LIMA Zhou et al.
-(2023) released a meticulously curated dataset comprising 1000 instructions, incorporating 750 popular queries and responses from community forums like Stack Exchange and wikiHow, alongside 250 expertly crafted entries.
-This compilation emphasizes the quality and diversity of instructions.
+(2023) aims to improve instruction-following ability of LLMs, consisting of 52K instructions, generated using OpenAI’s text-davincic-003.
+It has improved the self-instruction framework Wang et al.
+(2023b), leading to a greater diversity in the Alpaca dataset compared to the Self-Instruction dataset.
+Alpaca-cleaned yahma (2023) is a quality-enhanced version of the original Alpaca dataset, which addresses hallucinations, empty outputs, and incorrect answers, thereby improving the overall quality of the dataset.
+In addition, LIMA Zhou et al.
+(2023) has also released a carefully curated dataset, including 1000 instructions, which encompasses 750 popular queries and responses from community forums such as Stack Exchange and wikiHow, as well as 250 meticulously crafted entries.
 Instruction Fine-tuning Strategy The prevalent method of instruction fine-tuning, employed in models such as LIMA, Alpaca, Alpagasus and Superfiltering Li et al.
@@
 This approach, while emphasizing dataset quality and diversity Wang et al.
-(2023b); Li et al.
-(2023a), neglects the inherent complexity of the instruction sets.
+(2023b), neglects the inherent complexity of the instruction sets.
 Consequently, this one-off IFT fails to adequately equip models with the nuanced capability to comprehend and execute a wide array of instructions.
@@
 In contrast, within the Alpaca-clean dataset, instructions with a difficulty score lower than 1.5 account for 52.3%; in the LIMA dataset, the proportion of instructions with a difficulty score below 1.5 is merely 9.9%, while those scoring above 4.5 account for a significant 58.3%.
-Details of their comparison can be found in Appendix A.1.
+Details of their comparison can be found in Appendix A.1and A.2.
 This highlights the feasibility of employing GPT-4 to assess the difficulty level of instructions.
@@
 4 Experiments
-This paper seeks to validate the progressive alignment hypothesis, examining the relative efficacy of two instruction alignment training methodologies: One-off Instruction Fine-Tuning (One-off IFT) versus our proposed Phased Instruction Fine-Tuning (Phased IFT), through three distinct experiments.
-Firstly, Experiment 1 evidences Phased IFT’s superior performance over One-off IFT across two instruction data, establishing Phased IFT’s effectiveness.
-Secondly, Experiment 2 assesses the merit of stratifying instruction datasets by difficulty versus arbitrary segmentation, aiming to attribute Phased IFT’s success to difficulty-stratified instruction data categorization rather than to the phased training approach per se.
-Finally, Experiment 3, conducting a full permutation analysis of Phased IFT, seeks to demonstrate that the method’s success is specifically due to uptraining in a sequence of increasing difficulty, with alternative sequencing failing to produce comparable outcomes.
-Datasets and Evaluation We conducted extensive experiments on the standard Alpaca 52K instruction data and its refined version Alpaca-cleaned, with the following three recent LLMs: Llama-2 7/13B and Mistral-7B respectively, evaluating them across 6 benchmark datasets, including Self-Instruct of size 252, WizardLM of size 218, Koala of size 180, Vicuna of size 80, OASST of size 188 Köpf et al.
-(2023), and Anthropic Bai et al.
-(2022) of size 129.
-Following evaluation from Alpagasus Chen et al.
-(2024), we prompt GPT-4 to judge Win-Tie-Lose and compute a win rate.
-Under identical instruction and input, GPT-4 as an auto-grader Zheng et al.
+This paper aims to validate the progressive alignment hypothesis by comparing the efficacy of two instruction training methodologies: One-off Instruction Fine-Tuning (One-off IFT) and our proposed Phased Instruction Fine-Tuning (Phased IFT), through three experiments and ablation studies.
+Experiment 1 demonstrates the superior performance of Phased IFT over One-off IFT across two instruction datasets, establishing its effectiveness.
+Experiment 2 evaluates the effectiveness of stratifying instruction datasets by difficulty versus arbitrary segmentation, aiming to attribute Phased IFT’s success to difficulty-stratified instruction data categorization rather than to the phased training approach.
+Experiment 3 conducts a full permutation analysis of Phased IFT to demonstrate that its success is specifically due to uptraining in a sequence of increasing difficulty, with alternative sequencing failing to produce comparable outcomes.
+Datasets and Evaluation We conducted extensive experiments on the standard Alpaca 52K instruction data and its refined version, Alpaca-cleaned, using 6 recent LLMs: Llama-2 7/13/70B, Llama-3 8/70B and Mistral-7B.
+These models were evaluated across 6 benchmark datasets: Self-Instruction Wang et al.
+(2023b) of size 252, WizardLM Xu et al.
+(2024) 218, Koala Geng et al.
+(2023) 180, Vicuna Chiang et al.
+(2023) 80, OASST 188 Köpf et al.
+(2023), and Anthropic of size 129 Bai et al.
+(2022).
+Following the evaluation framework established by Alpagasus Chen et al.
+(2024), we prompted GPT-4 to assess Win-Tie-Lose outcomes and compute win rates across the aforementioned 6 benchmarks.
+Subsequently, a weighted average win rate metric is calculated based on these 6 benchmark datasets, referred to as Avg.
+win rate.
+Using identical instruction and input, GPT-4 acts as an auto-grader Zheng et al.
 (2023); Chia et al.
-(2023) to score the quality of output from both One-off IFT and Phased IFT methods, on a scale of 1 to 10 (See prompt in Appendix A.2).
-For fairness, we score each instruction twice, i.e., placing Phased IFT’s output before/after the One-off IFT’s output.
-An instruction’s overall outcome is defined by:
+(2023) to score the quality of output from both One-off IFT and Phased IFT methods on a scale of 1 to 10 (see prompt in Appendix A.3).
+To ensure fairness, each instruction is scored twice, with Phased IFT’s output positioned before and after One-off IFT’s output.
+An instruction’s overall outcome is determined based on these evaluations.
   • Win: win twice or win once and tie once;
@@
 Alpaca-cleaned 23.9 16.4 11.4 1.00 2.22 4.25 113.42 200.99 301.60
-Stratifying Alpaca by Difficulty Score
-Figure 4: Difficulty score histogram of Alpaca-cleaned data and its cumulative probability density.
-The instruction difficulty within the Alpaca and Alpaca-cleaned data are quantitatively assessed by GPT-4, assigning scores from 1 to 5, with higher score denoting increased complexity.
-Based on the density curve (e.g., Figure 2 for Alpaca and Figure 4 for Alpaca-cleaned), we adopt two threshold values, namely 1.5 and 3.5 for both Alpaca and Alpaca-cleaned, and segment them into three stages sub-datasets, respectively.
-We refer to as the Alpaca-3-stages and Alpaca-cleaned-3-stages.
-Additionally, to conduct ablation experiments, we also construct a randomly sampled 3-stages data, with each stage containing the same number of instructions as the difficulty-segmented subset.
-This is referred to as the Alpaca-rand-3-stages and Alpaca-cleaned-rand-3-stages data.
-Detailed statistics, as delineated in Table 1.
-Implementations We utilized Huggingface’s Trainer with DeepSpeed’s Zero3 for data-parallel, full-parameter fine-tuning in our training regimen.
-Consistently across experiments, a learning rate of 5e-6 and a duration of two training epochs were applied.
-We set a per-device batch size of 16 and a gradient accumulation steps of 1, a weight decay of 0.1, a warmup ratio of 0.1, and a cosine annealing learning rate scheduler.
-4.1 Experiment 1: Comparison win rate of Phased IFT and One-off IFT
-Table 2: Comparisons of win rates between Phased IFT and One-off IFT.
-Phased IFT involves up-training on Alpaca-3-stages and Alpaca-cleaned-3-stages, while One-off IFT is fine-tuned solely on the original Alpaca and Alpaca-cleaned data.
+Stratifying Alpaca by Difficulty Score The instruction difficulty within the Alpaca and Alpaca-cleaned datasets is quantitatively assessed by GPT-4, assigning scores from 1 to 5, with higher scores indicating increased complexity.
+Using density curves (e.g., Figure 2 for Alpaca and Figure 5 for Alpaca-cleaned), we adopt threshold values of 1.5 and 3.5 for both datasets, segmenting them into three stages sub-datasets, referred to as Alpaca-3-stages and Alpaca-cleaned-3-stages.
+For ablation experiments, we construct randomly sampled 3-stages datasets, with each stage containing the same number of instructions as the difficulty-segmented subset.
+These are denoted as Alpaca-rand-3-stages and Alpaca-cleaned-rand-3-stages.
+Detailed statistics are provided in Table 1.
+Implementations In our training, we employed Huggingface’s Trainer coupled with DeepSpeed’s Zero3 for data-parallel, full-parameter fine-tuning.
+Consistently, a learning rate of 5e-6 and a 2 epoch duration were utilized across experiments.
+We set a per-device batch size of 4 for 70B model and 16 for other models, a gradient accumulation step of 1, with a weight decay of 0.1 and a warmup ratio of 0.1.
+Additionally, a cosine annealing learning rate scheduler was employed.
+4.1 Experiment 1: compares the win rates of Phased IFT and One-off IFT
+Table 2: Comparisons of win rates of Phased IFT against One-off IFT.
 The changes in win rates are represented by percentage differences, where a "+" indicates an increase and "-" denotes a decrease in percentage.
-Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Data
-Llama-2 7B +5.75 +10.32 +4.17 +11.25 +8.78 +12.40 Alpaca
-Mistral-7B -0.20 -2.52 +6.67 +5.00 +3.19 +3.10
-Llama 2-13B +4.37 +6.88 +10.56 +11.25 +10.11 +13.18
-Llama 2-7B +3.77 +4.13 +4.17 +3.12 - - Alpaca-cleaned
-Llama 2-13B +2.78 +3.21 +1.39 +6.88 +7.45 +9.30
-Table 2 demonstrates that the Phased IFT method yields superior performance across a wide range of evaluation metrics compared to the One-off IFT method, with the sole exceptions being the Mistral-7B model’s lower win rates on the Self-Instruct and WizardLM benchmarks.
-Specifically, the Llama-2 7B model, when fine-tuned with the Phased IFT method on the Alpaca-3-stages instruction set, experienced a win rate increase ranging from a minimum of +4.17 absolute points (on Koala) to a maximum of +12.40 absolute points (on Anthropic).
-Similarly, the Llama-2 13B model’s win rate improved by at least +4.37 points (on Self-Instruct) to as much as +13.18 absolute points (on Anthropic).
-Notably, the Mistral-7B model saw a win rate increase of +6.67 absolute points on Koala and +3.10 points on Anthropic, but it exhibited a decrease in win rates of -0.20 and -2.52 absolute points on Self-Instruct and WizardLM, respectively.
-On the Alpaca-cleaned instruction set, the win rate of the Llama 2-7B model fine-tuned via Phased IFT improved by at least +3.12 absolute points (on Vicuna) and up to +4.17 points (on Koala), while the Llama-2 13B model saw increases ranging from +1.39 absolute points (on Koala) to +9.30 points on Anthropic.
-These results from Experiment 1 affirm the efficacy of the Phased IFT method for instruction-based fine-tuning of pre-trained LLMs, leading to notable performance enhancements on Llama-2 7B/13B and Mistral-7B models.
-4.2 Experiment 2: Comparison win rate of difficulty-stratified 3-stages and randomly sampled 3-stages
-Table 3: Comparison of win rates between Phased IFT and One-off IFT methods is outlined in three rows.
-The first row (Row #1) assesses Phased IFT on Alpaca-rand-3-stages against One-off IFT using the original Alpaca 52K data.
-In the second row (Row #2), the comparison is made between Phased IFT on Alpaca-cleaned-rand-3-stages and One-off IFT on the original Alpaca-cleaned data.
-The third row (Row #3) evaluates the win rates of Phased IFT on Alpaca-3-stages in contrast to Phased IFT on Alpaca-rand-3-stages.
-Each entry corresponds to Table 2.
-Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Row
-Llama-2 7B -1.19 -2.75 +0.56 +0.62 -0.66 +0.75 #1
-Llama-2 13B -3.97 -3.67 -3.89 -3.75 -2.93 -3.49
-Llama-2 7B -2.78 -2.52 -5.00 -1.88 - - #2
-Llama-2 7B +9.13 +12.16 +13.61 +10.62 +5.85 +6.59 #3
-Llama-2 13B +4.76 +7.57 +1.94 +10.00 +10.37 +10.85
-Table 3 provides insight into the effectiveness of our Phased IFT method for fine-tuning pre-trained models Llama-2 7B and 13B on randomly sampled 3-stages instruction data.
-Notably, the Phased IFT method, in comparison to the One-off IFT approach, exhibited a decline in win rates across several test sets.
-Specifically, as listed in the Row #1 fine-tuned with Phased IFT on Alpaca-rand-3-stages, the win rate of the Llama-2 13B model decreased significantly compared to One-off IFT on the original Alpaca 52K data, with reductions ranging from -2.93 absolute points on Anthropic to -3.97 points on Self-Instruct.
-The Row #2 of Table 3 which focuses on the application of Phased IFT for fine-tuning Llama-2 7B on Alpaca-cleaned-rand-3-stages, shows a pronounced decrease in win rate compared to the One-off IFT on the original Alpaca-cleaned data, with declines spanning from -1.88 absolute points on Vicuna to -5.00 absolute points on Koala.
-Contrastingly, Row #3 of Table 3 demonstrates that applying Phased IFT for uptraining the Llama-2 7B and 13B models on Alpaca-3-stages, as opposed to on Alpaca-rand-3-stages, yielded significant improvements in win rate.
-Specifically, the Llama-2 7B model saw increases from +5.85 absolute points on OASST to +13.61 on Koala, while the Llama-2 13B model’s performance improved by +1.94 points on Koala and up to +10.85 points on Anthropic.
-These results indicate that uptraining with randomly selected 3-stages instruction data does not inherently enhance performance and can lead to reductions across most test sets.
-This contrasts with experiment 1 where Phased IFT applied to difficulty-stratified 3-stages instruction sets, matched for instruction count and compared against the same baseline, yielded substantial performance gains.
-This highlights two key insights: firstly, the improvements observed are not solely attributable to the 3-stages uptraining process; secondly, categorizing instruction data by difficulty level provides a structured fine-tuning approach, enabling pre-trained models to progressively adapt to more complex tasks.
-Table 4: Win rate of all permutations of uptraining of Llama-2 7B/13B on Alpaca-3-stages instruction data, compared to One-off IFT on the original Alpaca data. Each entry is the same as in Table 2.
-Permutations Self-Instruct WizardLM Koala Vicuna OASST Anthropic Base Models
-1-2-3 +5.75 +10.32 +4.17 +11.25 +8.78 +12.40 7B
-2-1-3 +0.40 +2.98 +0.56 +11.25 +6.65 +8.14
-3-1-2 +3.57 +1.15 -3.33 +12.50 +1.06 +5.43
-1-3-2 -0.79 -2.52 -0.83 +2.50 -2.39 +5.43
-2-3-1 -5.16 -9.40 -13.61 -8.12 -5.32 -5.04
-3-2-1 -4.76 -5.96 -9.17 -6.25 -9.04 -1.16
-1-2-3 +4.37 +6.88 +10.56 +11.25 +10.11 +13.18 13B
-2-1-3 +6.15 +7.11 +10.28 +5.62 +5.85 +8.53
-3-1-2 +0.79 +1.38 +4.44 -6.88 +2.39 +2.33
-1-3-2 +1.39 +3.90 +5.00 +1.25 +1.06 +6.98
-2-3-1 -3.37 -6.65 -1.11 -7.50 -11.70 -10.08
-3-2-1 -5.16 -5.28 -5.00 -11.88 -12.77 -7.36
-4.3 Experiment 3: Comparison win rate of uptraining on Alpaca-3-stages in all permutations
-To investigate the impact of varying the sequence of Alpaca-3-stages on the performance of the Llama-2 7B/13B models with clarity and precision, the following structure is proposed.
+Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Avg. win-rate Data
+Llama-2 7B +4.17 +6.42 +5.56 +6.88 +9.31 +14.34 +7.26 Alpaca
+Mistral 7B -0.60 +5.96 +6.11 +16.25 +7.98 +12.02 +6.30
+Llama-2 13B +1.98 +2.75 +7.78 +11.88 +11.44 +16.28 +7.35
+Llama-2 70B -3.04 +2.57 +7.27 +13.46 +4.10 +0.00 +2.82
+Llama-3 8B +3.05 +9.67 +9.52 +7.50 +6.32 +12.60 +7.64
+Llama-3 70B +1.65 +10.10 +5.23 +5.84 +2.34 +7.87 +5.23
+Llama-2 7B +4.76 +1.15 +4.44 +2.50 +5.85 +1.16 +3.53 Alpaca- cleaned
+Llama-2 13B +4.56 +2.98 +5.83 +15.62 +6.91 +10.85 +6.49
+Llama-2 70B +0.40 +6.42 +8.33 +12.50 +10.90 +14.73 +7.59
+Llama-3 8B +5.75 +3.90 +1.15 +0.62 +4.79 +5.43 +3.97
+Table 2 lists the win rate results of Phased IFT compared to the One-off IFT method, demonstrating that our proposed Phased IFT significantly outperforms the One-off IFT method across six benchmarks on two training datasets.
+The first part of Table 2 compares the results of the two fine-tuning methods on Alpaca.
+Phased IFT was uptrained on Alpaca-3-stages, while One-off IFT was trained once on the original Alpaca.
+Using six base models, Phased IFT exceeded One-off IFT in the Avg.
+win-rate metric.
+Specifically, compared to the One-off IFT method trained once on the original Alpaca dataset, the Llama-2 7B using the Phased IFT fine-tuning method on Alpaca-3-stages showed a minimum win rate increase of +4.17 on Self-Instruct and a maximum of +14.34 on Anthropic, with an average win rate of +7.26.
+The Mistral 7B also achieved an Avg.
+win rate of +6.30; Llama 2 13B reached an average win rate of +7.35, with a win rate of +16.28 on Anthropic; Llama 3 8B and 70B both achieved an average win rate increase of over +5.0.
+The second part of Table 2 compares the win rates of five base models using the two fine-tuning methods on Alpaca-cleaned.
+Specifically, the three Llama2 series models achieved average win rates of +3.53, +6.49, and +7.59, respectively, while Llama3 8B achieved an average win rate of +3.97.
+It can be observed that the average win rates of Llama2 7B/13B and Llama3 8B on Alpaca-cleaned are lower than their results on Alpaca, while Llama2 70B’s average win rate increased from +2.82 on Alpaca to +7.59 on Alpaca-cleaned.
+Due to time constraints, experiments on Llama3 70B have not yet been completed.
+The results of Experiment 1 indicate that the Phased IFT fine-tuning method is indeed more effective than One-off IFT, achieving significantly higher win rates on most of the six benchmarks, with an overall average win rate significantly higher than that of One-off IFT.
+4.2 Experiment 2: Comparison of win rate between difficulty-stratified 3-stages and randomly sampled 3-stages
+Table 3: Part #1 assesses win rates of Phased IFT on Alpaca-rand-3-stages against One-off IFT on the original Alpaca data.
+Part #2 evaluates win rates of Phased IFT on Alpaca-cleaned-rand-3-stages over One-off IFT on the original Alpaca-cleaned data.
+Part #3 measures win rates of Phased IFT on Alpaca-3-stages compared to Phased IFT on Alpaca-rand-3-stages.
+Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Avg. win-rate Part
+Llama-2 7B -0.60 +2.98 -6.61 +3.12 -1.33 +1.94 -0.42 #1
+Llama-2 13B -1.98 -2.29 -0.29 -6.88 -2.93 -0.39 -2.10
+Llama-2 70B -10.32 -5.96 -6.90 -3.12 -11.97 -5.43 -7.96
+Llama-3 8B -7.74 -1.38 -1.15 +5.00 +0.00 -2.33 -2.25
+Llama-3 70B -10.00 -3.63 -4.44 -5.63 -6.89 -2.73 -5.92
+Llama-2 7B -1.34 -5.07 -3.85 +8.12 -1.34 -2.76 -2.00 #2
+Llama-2 13B +0.58 -0.69 +1.26 +1.88 +2.97 -0.39 0.84
+Llama-3 8B +2.80 +0.69 -6.94 +2.50 +4.30 -6.59 -0.22
+Llama-2 7B +2.38 +3.90 +11.78 +11.25 +2.39 +8.53 +5.74 #3
+Llama-2 13B +6.94 +2.98 +6.32 +16.25 +11.70 +12.79 +8.29
+Llama-2 70B +8.93 +5.50 +10.34 +19.38 +11.97 +8.14 +9.71
+Llama-3 8B +2.18 +0.69 +13.79 +2.50 +8.78 +9.30 +5.95
+Llama-3 70B +8.07 +14.10 +12.50 +17.12 +8.63 +5.96 +10.62
+Table 4: Win rates of all permutations of Phased IFT compared to One-off IFT on the original Alpaca data.
+Permutations Self-Instruct WizardLM Koala Vicuna OASST Anthropic Avg. win-rate Base Models
+1-2-3 +4.17 +6.42 +5.56 +6.88 +9.31 +14.34 +7.26 Llama2 7B
+2-1-3 +1.39 +6.65 +2.30 +6.88 +4.79 +8.91 +4.59
+3-1-2 +0.80 +0.94 -2.35 +6.88 +3.83 +8.66 +2.26
+1-3-2 +3.57 -1.15 -1.15 +1.88 +2.66 +8.91 +2.14
+2-3-1 -2.61 -6.81 -7.65 -5.70 -4.12 -0.39 -4.5846
+3-2-1 -7.23 -3.18 -9.09 +1.54 -5.98 +2.73 -4.5847
+1-2-3 +1.98 +2.75 +7.78 +11.88 +11.44 +16.28 +7.35 Llama2 13B
+2-1-3 +2.33 +1.63 +8.82 +6.39 +6.41 +12.29 +5.57
+3-1-2 -0.61 +0.70 +4.39 +5.62 +2.75 +8.40 +2.71
+1-3-2 -0.40 +2.14 +6.17 +1.25 +4.72 +9.60 +3.53
+2-3-1 -6.83 -10.56 -8.33 -8.12 -3.57 +0.78 -6.44
+3-2-1 -5.42 -8.22 -7.69 -8.23 -10.50 -1.97 -7.09
+1-2-3 +3.05 +9.67 +9.52 +7.50 +6.32 +12.60 +7.64 Llama3 8B
+2-1-3 -2.48 +3.61 +4.94 +3.38 +6.21 +5.24 +3.02
+3-1-2 +0.82 +2.12 +0.90 +3.80 +2.53 +6.35 +2.32
+1-3-2 -0.41 +0.00 +0.31 -1.97 +5.40 +5.16 +1.40
+2-3-1 -11.27 -14.44 -11.70 -14.86 -16.44 -14.22 -13.57
+3-2-1 -11.88 -9.39 -5.00 -5.77 -9.23 -3.97 -8.26
+Table 5: Ablation studies of various methods on Alpaca-3-stages instruction data. Each entry denotes the win rate of Phased IFT against One-off IFT on the original Alpaca data.
+Method Self-Instruct WizardLM Koala Vicuna OASST Anthropic Avg. win-rate Models
+Phased IFT stage1 -8.53 -10.32 -15.00 -6.25 -10.11 -4.26 -9.59 Llama2 7B
+stage2 +3.57 +2.98 -0.83 +5.00 +3.99 +7.36 +3.34
+stage3 +4.17 +6.42 +5.56 +6.88 +9.31 +14.34 +7.26
+IFT on stage2 -0.20 +2.52 -0.83 +8.75 +0.53 +0.39 +1.14
+IFT on stage3 +0.99 +6.88 +1.67 +16.25 +0.00 +10.08 +4.44
+IFT on mixed stage1&2 +2.78 +1.83 -8.33 -5.00 -4.79 +0.00 -1.62
+IFT on mixed stage1&3 -2.58 +0.00 -2.50 +3.12 -3.46 -3.88 -1.91
+IFT on mixed stage2&3 +0.40 +1.38 +4.17 +8.12 +1.33 +12.79 +3.53
+Phased IFT stage1 -6.55 -10.09 -13.61 -6.88 -6.65 -0.78 -7.83 Llama2 13B
+stage2 +0.40 -0.23 +4.44 +3.75 +3.72 +16.28 +3.77
+stage3 +1.98 +2.75 +7.78 +11.88 +11.44 +16.28 +7.35
+IFT on stage2 -2.40 -0.47 +4.09 +3.80 +6.87 +9.06 +2.66
+IFT on stage3 +4.56 +7.26 +4.11 +10.12 +3.26 +9.18 +5.80
+IFT on mixed stage1&2 -0.60 -1.61 -3.06 +0.62 +2.93 +2.71 -0.09
+IFT on mixed stage1&3 +1.39 +0.46 -3.33 +0.00 +6.12 +3.10 +1.33
+IFT on mixed stage2&3 +1.39 +3.67 +7.50 +8.12 +7.45 +10.85 +5.68
+Experiment 1 demonstrates the effectiveness of Phased IFT compared to One-off IFT, while Experiment 2 reveals that this effectiveness arises from multi-stage sub-datasets with increasing difficulty, not from the multi-stage training itself.
+Therefore, we designed multi-stage datasets with random segmentation, each segment maintaining the same quantity as the corresponding dataset with increasing difficulty, but the data points were randomly sampled.
+Table 3 lists the results of three sets of experiments as follows.
+The first part of Table 3 presents the win rate of Phased IFT on Alpaca-rand-3-stages relative to One-off IFT on the original Alpaca.
+Specifically, we used five baseline models, namely Llama2 7/13/70B and Llama3 8/70B.
+Their performance on six benchmarks was consistently lower than that of One-off IFT, indicated by negative win rates.
+This suggests that segmenting the dataset randomly into multiple subsets and then performing uptraining did not yield benefits.
+Similarly, the second part of Table 3 shows the win rate of Phased IFT on Alpaca-cleaned-rand-3-stages relative to One-off IFT on the original Alpaca-cleaned.
+Based on the results, the win rates of Llama2 7B and Llama3 8B were negative, indicating that multi-stage training did not help.
+The third part of Table 3 shows the win rate of Phased IFT on Alpaca-3-stages relative to Phased IFT on Alpaca-rand-3-stages.
+The results indicate that uptraining on the difficulty-increasing Alpaca-3-stages dataset resulted in significantly higher win rates than uptraining on the randomly segmented Alpaca-rand-3-stages dataset.
+Specifically, we used five baseline models for uptraining on both Alpaca-3-stages and Alpaca-rand-3-stages.
+The average win rate of Llama3 70B reached +10.62, Llama2 70B reached +9.71, Llama2 13B reached +8.29, and Llama2 7B and Llama3 8B also reached +5.74 and +5.95, respectively.
+These models all achieved consistently high win rates, far higher than uptraining on randomly segmented subsets.
+These experiments indicate that uptraining on multi-stage datasets with increasing difficulty is beneficial and enhances the model’s capabilities.
+Uptraining on randomly segmented multi-stage datasets reduced the model’s capabilities, and the results were not as good as single-step fine-tuning on the original dataset (For a detailed graphical presentation, please refer to Appendix A.4).
+This suggests that multi-stage training methods did not improve the model’s capabilities and yield benefits.
+The effect of Phased IFT mainly comes from multi-stage subsets with increasing difficulty, which also validates our proposed progressive alignment hypothesis, that aligning pre-trained models to accomplish the final human task is a gradual process.
+4.3 Experiment 3: compares the win rates achieved by uptraining on Alpaca-3-stages across all permutations
+To investigate the impact of varying the sequence of Alpaca-3-stages on the performance of the Llama-2 7B/13B and Llama3 8B models with clarity and precision, the following structure is proposed.
 This structure presents the uptraining order across the three stages, denoted as Stage 1 (1), Stage 2 (2), and Stage 3 (3), to illustrate the sequence in which the model was uptrained on different segments of the instruction data:
@@
   • 3-2-1: This order shows that the model was first uptrained on Stage 3 data, then Stage 2, and finally Stage 1.
-Table 4 demonstrates the performance enhancements achieved by fine-tuning the pretrained Llama-2 7B/13B models through the Phased IFT method across 6 permutations of these stages, in contrast to the One-off IFT method applied directly on the original Alpaca data.
-The results distinctly reveal that the Llama-2 7B/13B models achieve maximal win-rate improvements when sequentially fine-tune in the 1-2-3 order using Phased IFT, substantially outperforming other sequential arrangements.
-The Llama-2 7B model, after undergoing Phased IFT in a 1-2-3 sequence, demonstrates significant win-rate enhancements of +5.75, +10.32, +4.17, +11.25, +8.78, and +12.40 absolute points across Self-Instruct, WizardLM, Koala, Vicuna, OASST, and Anthropic benchmarks, respectively.
-This outperformance is substantial when compared to the model fine-tuned via the One-off IFT method on the 52K Alpaca data, as well as models fine-tuned in alternative sequences.
-Notably, the 1-2-3 and 2-1-3 training sequences yield markedly superior win rate than the 3-1-2 and 1-3-2 sequences, which in turn outperform the 2-3-1 and 3-2-1 sequences.
-The latter sequences exhibit win rate significantly lower than those achieved through One-off IFT.
-Similarly, the Llama-2 13B model, fine-tuned in the 1-2-3 sequence using Phased IFT, show win-rate improvements of +4.37, +6.88, +10.56, +11.25, +10.11, and +13.18 absolute points on the same benchmarks, respectively, again surpassing the performance of models fine-tuned through other sequences or the One-off IFT method.
-The training sequences of 1-2-3 and 2-1-3 consistently resulted in higher win rate compared to 3-1-2 and 1-3-2, which outperform 2-3-1 and 3-2-1 sequences.
-The win rate for the 2-3-1 and 3-2-1 sequences are notably inferior to those achieved with One-off IFT.
-These findings compellingly support the hypothesis that uptraining in order of increasing difficulty presents significant positive gains, whereas reverse order training yields negligible benefits.
-This underpins our progressive alignment hypothesis, positing that the process of aligning pretrained models to human intent through supervised instruction-based fine-tuning is inherently gradual.
-Beginning with basic and straightforward instructions before progressively tackling more challenging ones ensures a systematic alignment with end tasks.
-This approach not only validates the utility of constructing diverse and difficulty-tiered instruction data but also highlights the strategic importance of adopting a progressively challenging fine-tuning regimen to enhance model performance and alignment with human intent in task completion.
-To summarize, Experiment 1 validates the superiority of phased instruction fine-tuning (Phased IFT) over one-off IFT, evidenced by a notable enhancement in win rates.
-Experiment 2 underscores the significance of stratifying instruction dataset by difficulty score, yielding a pronounced improvement in win rate against random sampling for uptraining fine-tuning, as depicted in Row #3 of Table 3, with the former method achieving a significant uplift in win rate.
-Experiment 3 demonstrates that sequential uptraining fine-tuning by ascending difficulty represents the optimal phased training approach, concurrently corroborating the progressive alignment hypothesis.
-Collectively, these experiments show the efficacy of the progressive alignment hypothesis and Phased IFT.
+Table 4 presents the uptraining results of Llama2 7/13B and Llama3 8B models on the three-stages Alpaca-3-stages dataset, evaluating 6 possible permutations.
+Analyzing the average win rate, a clear pattern emerges: the sequences 1-2-3 and 2-1-3 outperform 1-3-2 and 3-1-2, which in turn outperform 2-3-1 and 3-2-1.
+This suggests that uptraining on sub-datasets of increasing difficulty can indeed enhance the model’s potential.
+The first part of Table 4 details the full permutation experiments on the Llama2 7B model.
+The 1-2-3 sequence achieved the highest win rates across 6 benchmarks and the average win rate, except for a slight dip below the 2-1-3 sequence in the WizardLM benchmark (e.g., +6.42 vs.
++6.65).
+The average win rate for the 1-2-3 sequence was +7.26, significantly higher than the other permutations.
+The 2-1-3 sequence also performed well with an average win rate of +4.59, both sequences sharing a common feature of placing the most difficult sub-dataset (stage 3) at the final stage of uptraining.
+The sequences 3-1-2 and 1-3-2, which place the medium difficulty sub-dataset (stage 2) last, also showed positive win rates of +2.26 and +2.14, respectively.
+Surprisingly, when the easiest sub-dataset (stage 1) was placed last, the sequences 2-3-1 and 3-2-1 had lower average win rates than One-off IFT on the original Alpaca dataset.
+The second part of Table 4 shows the full permutation experiments on the Llama2 13B model.
+Similarly, the 1-2-3 and 2-1-3 sequences achieved the highest average win rates of +7.35 and +5.57, respectively, significantly outperforming the other permutations.
+These sequences also shared the characteristic of placing the most difficult sub-dataset (stage 3) last.
+The 1-3-2 and 3-1-2 sequences achieved average win rates of +3.53 and +2.71, respectively, with the 3-2-1 sequence having the lowest average win rate of -7.09.
+Interestingly, the permutation experiments on the Llama2 13B model aligned perfectly with the win rate predictions of the progressive alignment hypothesis, ranking the uptraining results from highest to lowest average win rate as follows: 1-2-3 > 2-1-3 > 1-3-2 > 3-1-2 > 2-3-1 > 3-2-1.
+This further confirms the benefit of uptraining on sub-datasets of increasing difficulty when fine-tuning LLMs.
+The third part of Table 4 presents the full permutation experiments on the Llama3 8B model.
+The 1-2-3 and 2-1-3 sequences achieved the highest average win rates of +7.64 and +3.02, respectively, outperforming the 3-1-2 and 1-3-2 sequences, which had average win rates of +2.32 and +1.40, respectively.
+Unlike the Llama2 7/13B models, the 2-3-1 sequence had the lowest win rate at -13.57, significantly lower than the -8.26 of the 3-2-1 sequence.
+Further investigation is needed to understand the reasons behind this result.
+In sum, the results indicate a phenomenon where placing more difficult sub-datasets at the final stage of uptraining on differentiated difficulty sub-datasets can significantly enhance the model’s capabilities.
+This approach markedly increases the model’s win rate, making the generated results closer to human perception, as the judgments of GPT-4 align with human approximately 75%.
+4.4 Ablation Studies
+We conducted ablation experiments using Llama2 7B/13B on Alpaca-3-stages, incorporating various sub-dataset combinations.
+These combinations included performing Phased IFT on Alpaca-3-stages, IFT exclusively on stage 2, IFT exclusively on stage 3, IFT on a combined sub-dataset of stage 1 and stage 2, IFT on a combined sub-dataset of stage 1 and stage 3, and IFT on a combined sub-dataset of stage 2 and stage 3.
+Table 5 presents two sets of ablation experiments, further verifying the role of each sub-dataset in fine-tuning LLMs.
+The first part of Table 5 shows the ablation results for the Llama2 7B base model.
+Performing IFT exclusively on stage 2 yielded a positive gain, with a win rate of +1.14.
+The win rate for Phased IFT at the stage 2 phase was +3.34, indicating that stage 1 also played a positive role during uptraining.
+The win rate for IFT exclusively on stage 3 reached +4.44, which was higher than the Phased IFT at stage 2 but lower than the Phased IFT at stage 3.
+This suggests that stage 3 played a major role in performance improvement during Phased IFT.
+This result is consistent with the conclusions of Alpagasus Chen et al.
+(2024), which indicate that high-quality data filtered by GPT is crucial for fine-tuning models.
+Performing IFT on the combined sub-dataset of stage 1 and stage 2 resulted in a negative win rate, as did performing IFT on the combined sub-dataset of stage 1 and stage 3.
+This implies that redundant datasets are not conducive to instruction training.
+However, performing IFT on the combined sub-dataset of stage 2 and stage 3 resulted in a win rate of +3.53, higher than the +1.14 for IFT on stage 2 alone but lower than the +4.44 for IFT on stage 3 alone.
+This nuanced result suggests that mixing stage 2 data had a detrimental effect, leading to reduced model performance.
+Nevertheless, when uptraining across all three stages (stage 1, stage 2, and stage 3), the win rate steadily improved, indicating a gradual enhancement in model capability.
+This is a very interesting experimental phenomenon, and similar observations were made in ablation experiments with Llama2 13B.
+This indicates that the model performance improvement achieved through uptraining on increasingly difficult sub-datasets is not a trivial training technique.
+The quality of the instruction data and the alignment method are key factors in fine-tuning of large language models.
+This also further supports the hypothesis of alignment progression as a genuine phenomenon.
 5 Conclusion
-The paper introduces the progressive alignment hypothesis, suggesting that the transition of a pretrained model’s next-word prediction capability into an ability to follow human intentions is a sequential learning process.
-This process entails progressively aligning with human intent, advancing from simpler to more complex tasks.
-In light of this hypothesis, we propose a phased instruction fine-tuning (Phased IFT) approach.
-Utilizing GPT-4 to score the difficulty of instructions, we categorize the instruction dataset into different stages according to the difficulty score density distribution.
-Diverging from prior methods that apply a one-off instruction fine-tuning (One-off IFT) on the whole instruction data, our approach entails staged, supervised uptraining, addressing each stage in ascending order of difficulty.
-Through extensive experiments, we have validated the progressive alignment hypothesis and the efficacy of Phased IFT, demonstrating its utility in developing diverse instruction dataset and enhancing instruction fine-tuning.
+The paper introduces the progressive alignment hypothesis, and then proposes Phased IFT approach.
+Utilizing GPT-4 to score the difficulty of instructions, the instruction dataset is categorized into different stages according to the difficulty.
+Diverging from prior methods that apply a One-off IFT on the whole instruction data, our approach entails staged, supervised uptraining on each stage in ascending order of difficulty.
+Through extensive experiments, we have validated the progressive alignment hypothesis and the efficacy of Phased IFT, demonstrating its utility in developing diverse instruction dataset and enhancing fine-tuning LLMs.
 6 Broader Impact
-This paper discusses a fundamental issue in Large Language Models (LLMs): how to align LLMs more efficiently.
-The work introduces a simple and effective method for alignment, which involves analyzing the difficulty of instruction dataset and structuring the dataset into several stages, followed by supervised fine-tuning through uptraining.
-We conduct three experiments, with an additional experiment detailed in Appendix A.3, to actively validate the existence of the progressive alignment hypothesis.
-If the progressive alignment hypothesis is a pervasive phenomenon, this work is poised to have a widespread impact on the field of LLMs.
+This paper discusses a fundamental problem in Large Language Models (LLMs): How to more effectively utilize instruction datasets to align LLMs more efficiently?
+This paper introduces a simple and effective alignment method, which includes the analysis of instruction difficulty and structuring the dataset into multi-stage sub-datasets, with each sub-dataset increasing in difficulty.
+Supervised fine-tuning is then carried out through uptraining.
+We conducted a series of experiments to actively verify the existence of the progressive alignment hypothesis.
+If the progressive alignment hypothesis is a universal phenomenon, then this work will have a broad impact on the direction of fine-tuning instructions for Large Language Models.
 The datasets employed in this work, namely Alpaca 52K and Alpaca-cleaned, are publicly available, and our training framework leverages open-source tools, utilizing Hugging Face’s Trainer and DeepSpeed.
-Furthermore, we have purchased access to OpenAI’s API to employ GPT-4 for assessing the difficulty of instruction datasets and for evaluating the quality of generated responses.
-We will make all our code and data openly available, including training code, GPT-4 difficulty scoring code, win-rate calculation code, and the difficulty-stratified Alpaca and Alpaca-cleaned instruction datasets.
+Furthermore, we have purchased access to OpenAI’s API to employ GPT-4 for assessing the difficulty of instructions and for evaluating the quality of generated responses.
+We make all our code and data openly available at https://github.com/xubuvd/PhasedSFT, including training codes, GPT-4 difficulty scoring codes, win-rate calculation codes, and the difficulty-stratified Alpaca and Alpaca-cleaned instruction datasets.
 7 Limitations
 Here, we discuss some limitations of our study to inspire future research in this direction.
-Automatic instruction dataset segmentation This work employs heuristic information from the cumulative density curve of difficulty scores, supplemented by human judgment for threshold selection.
-A significant challenge lies in developing algorithms capable of autonomously identifying threshold points from this curve, eliminating the reliance on human intuition and ensuring balanced data size across multi-stages sub-datasets.
-Validation on larger-scale LLMs Our empirical analysis is confined to pre-trained models with capacities of 7B and 13B, lacking empirical validation on larger-scale models, such as 65B or 70B.
-This limitation suggests a necessity for extending validation efforts to encompass a wider range of model sizes to enhance the generalizability of our findings.
+Automatic instruction dataset segmentation This study draws heuristic information from the cumulative density curve of difficulty scores in the instruction dataset using human experience, selecting several thresholds to divide the instruction dataset into multiple subsets at different stages, with increasing difficulty levels.
+Utilizing human experience to select thresholds is a significant task that not only involves partitioning the dataset into subsets from easy to difficult but also requires considering the balance in the quantity of each subset, heavily relying on human expertise.
+Therefore, a significant challenge lies in developing an algorithm capable of autonomously identifying threshold points from this cumulative density curve, eliminating the dependence on human intuition, and ensuring a balanced quantity across the multi-stage subsets
 Acknowledgements
@@
 Appendix A Appendix
-A.1 Histogram and density curve for Alpaca, Alpaca-cleaned and LIMA
+Table 6: Comparison of percentage distribution of difficulty scores on the Alpaca52K dataset with GPT-3.5 and GPT-4
+GPT [1.0, 2.5) [2.5, 3.5] (3.5, 4.5) [4.5, 5.0]
+3.5 2.18% 62.28% 25.96% 9.56%
+4 76.3% 13.38% 4.1% 6.17%
 Figure 4: Histogram and cumulative probability density of difficulty scores for Alpaca 52K dataset.
@@
 Figure 6: Histogram and cumulative probability density of difficulty scores for LIMA 1000 dataset.
+Figure 7: A prompt to ChatGPT-4 for rating model’s responses with a score between 1 and 10.
+Figure 8: Comparison of the win rate trends for uptraining on multi-stage sub-datasets with increasing difficulty (solid line) and uptraining on randomly divided multi-stage sub-datasets (dashed line).
+The gray horizontal line in the middle represents the baseline of One-off IFT on the original dataset.
+A.1 Comparison of percentage distribution of difficulty scores on the Alpaca52K dataset using GPT-3.5 and GPT-4
+In our initial assessment, we compared the differences of GPT-3.5 and GPT-4 (gpt-4-0613) in assigning difficulty scores to the Alpaca 52K dataset through human evaluation, the score ranges from 1.0 to 5.0.
+We divided the scores between 1.0 and 5.0 into four intervals, with the quantity percentage of each interval shown in the Table 6.
+The distribution of scores reveal a noticeable regression toward the mean for GPT-3.5, with scores between 2.5 and 3.5 accounting for 62.28% of the total.
+A manual review of 200 samples indicated a mismatch between the assigned difficulty levels and human judgment, particularly noting an excess of simpler questions within the dataset.
+In contrast, GPT-4’s difficulty ratings aligned with human evaluations for approximately 75% of the cases, with scores ranging between 2.5 and 3.5 comprising 13.38% of the total.
+A.2 Histogram and density curve for Alpaca, Alpaca-cleaned and LIMA
 We conducted a comparative analysis on the instruction difficulty of three datasets—Alpaca, Alpaca-cleaned, and LIMA—employing GPT-4 for scoring.
-Figures 5, 6 to 7 illustrate their respective score distributions and probability density functions.
+Figures 4, 5 to 6 illustrate their respective score distributions and probability density functions.
 Alpaca, generated by the Self-Instruct framework, consists of 52K triplets (instruction, input, output).
@@
 The logo should also contain an image of a fork and spoon in the same color.
-A.2 ChatGPT-4 for evaluating win rate
-Figure 8: A prompt to ChatGPT-4 for rating model’s responses with a score between 1 and 10.
+A.3 ChatGPT-4 for evaluating win rate
 Figure 7 details the prompt used for scoring and comparing the quality of output generated by two models using GPT-4.
@@
 Specifically, we provide GPT-4 with inputs in two different orders: (instruction, input, output-1 of model 1, output-2 of model 2) and (instruction, input, output-2 of model 2, output-1 of model 1) for separate evaluations.
-A.3 Experiment 4: Comparison win rate of IFT only on the Alpaca stage3 data vs. IFT on all the 52k Alpaca
-Table 5: Win rate comparison of One-off IFT only on the third stage selected from Alpaca-3-stages against One-off IFT on the original Alpaca 52K data, using Llama-2 7B model.
-Each entry is the same as in Table 2.
-Self-Instruct WizardLM Koala Vicuna OASST Anthropic
-+0.79 +6.65 +2.50 +5.00 +3.81 +4.75
-We compared the win rates of IFT conducted solely on the third stage, which includes 6.8K samples, against IFT performed on the entire 52K Alpaca data, as shown in Table 5.
-It was observed that the win rate for IFT on just stage 3 exceeds that of the full 52K dataset, aligning with conclusions drawn in Alpagasus Chen et al.
-(2024) that the quality of the instruction dataset is crucial.
-However, a comparison with Experiment 3 in Table 4 reveals that the win rate for IFT solely on stage 3 is significantly lower than when Phased IFT is conducted across all three stages, roughly only half of the latter’s.
-This underscores the existence of the progressive alignment hypothesis and validates the effectiveness of the Phased IFT approach.
+A.4 Comparison of win rate trends across multi-stages of uptraining
+Figure 8 is a summary of this paper, which incorporates the results from Table 3.
+As can be seen from the figure, with the progression of uptraining, it demonstrates the winning rate growth trend of five LLMs on multi-stage sub-datasets with increasing difficulty.
+This forms a stark contrast to the winning rate trend of the same five LLMs on multi-stage sub-datasets with randomly distributed difficulty levels.
+To be specific, the winning rate steadily continues to increase when uptraining on the sub-datasets with incremental difficulty.
+Conversely, when uptraining on the sub-datasets with random difficulty, the winning rate first increases then decreases, apart from the continuously declining winning rate trend of Llama3 70B.
+This indicates that the Progressive Alignment Hypothesis we proposed is an effective instruction learning method.
+It gradually learns from the pre-training model’s ability to predict the next word to eventually acquire the ability to complete the human end task.